OpenStudy (anonymous):

please explain y=(11/3)cos^-1(5-2x)

OpenStudy (anonymous):

I assume you mean: \[y=\frac{11cos^{-1}(5-2x)}{3}\] \[\text{ Where }cos^{-1}\text{ is the inverse cosine function}\]

OpenStudy (anonymous):

yes that is correct. I know \[\cos^1=-1/\sqrt{1-(5-2x^2) but don't know what to do after that

OpenStudy (anonymous):

yes that is correct. I know \[\cos^1=-1/\sqrt{1-(5-2x^2) but don't know what to do after that